OpenStudy (anonymous):

x²-6x-8=0 solve by completing the square

OpenStudy (bahrom7893):

x^2-6x-8+17=17

OpenStudy (anonymous):

\[x^2-6x=8\] \[(x-3)^2=8+9=17\] \[x-3=\pm\sqrt{17}\] \[x=3\pm\sqrt{17}\]

OpenStudy (anonymous):

satellite can u tell me how the heck ur typing like that

OpenStudy (anonymous):

you mean why or how?

OpenStudy (anonymous):

the "how" is latex